OpenStudy (swag):

80c=2L+8w What is the Length??? Someone show me how to solve please.

terenzreignz (terenzreignz):

Are we looking for a numerical solution? Or just solve for L?

OpenStudy (phi):

you want L by itself. the first step is add -8w to both sides of the equation what do you get ?

OpenStudy (phi):

just write -8w on both sides

OpenStudy (swag):

ok

OpenStudy (phi):

what did you get ?

OpenStudy (phi):

it should look like this 80c -8w =2L+8w - 8w now simplify the right side. you have 8 w's take away 8 w's. how many w's on the right side ?

OpenStudy (swag):

Sorry internet keeps going out. Anyway how many w's on the right??? Theres 2 on the right.

OpenStudy (phi):

t should look like this 80c -8w =2L+8w - 8w how many w's on the right ? think like this I have 8 w's and take away 8 w's

OpenStudy (swag):

Lol wouldn't that leave me with 0 if I take away 8 from 8. Sorry if I seem slow, that is just what I got.

OpenStudy (phi):

yes, zero w's (which is why we added -8w in the first place) you now have after simplifying 80c -8w =2L+8w - 8w 80c -8w =2L or 2L = 80c -8w now if 2 L's are 80c -8w one L will be have of 80c-8w divide both sides by 2: \[ \frac{2L}{2}= \frac{80c -8w}{2} \]

OpenStudy (phi):

*half (not have)

OpenStudy (phi):

remember these rules: if you have the same number in the top and bottom, they divide out (make 1) so we can simplify also \[ \frac{A+B}{2} = \frac{A}{2} + \frac{B}{2} \]

OpenStudy (anonymous):

L=40c-4w